vensania:

.) A sample containing both NaCl and NaBr, weighs 0.6400 g and gives a precipitate of AgCl and AgBr that weighs 0.5269 g. Another 0.6400 g sample was titrated with 0.1084 M AgNO3requiring 27.52 mL. Calculate the % w/w of NaCl and NaBr in the sample.

lilianamendez2:

@ax

Walen:

W(NaCl+NaBr+ inert substance ) = 1g Let Wt. of NaCl is xg & Wt. of NaBr is yg (NaCl+NaBr)+AgNO 3 ​ →(AgCl+AgBr)↓ xg yg excess 0.526g 58.5 x ​ 103 y ​ Cl 2 ​ ↓△ AgCl(0.426g) 58.5 x ​ moles of NaCl will produce 58.5 x ​ moles of AgCl ∴ ........................................................... 58.5 x ​ ×143.5g 103 y ​ mole s of NaBr will produce 103 y ​ mole of AgBr ∴ ........................................................... 103 y ​ ×183g A/q, 58.5 x ​ ×143.5+ 103 y ​ ×183=0.526g


(1) Now when Cl 2 ​ gas is passed over (AgCl+AgBr) mixture then AgBr will convert into AgCl AgBr+Cl 2 ​ →AgCl 103 y ​ moles → 103 y ​ moles ∴Wt. of AgCl=0.426g produce from AgBr= 103 y ​ ×143.5g ∴ A/q , Total Wt. of AgCl=0.426g ∴ 58.5 x ​ ×143.5+ 103 y ​ ×143.5=0.426g
(2) Solving (1) & (2) , we have x=0.0425g y=0.232g ∴ % of NaCl= 1 x ​ ×100=4.25 % % of NaBr= 1 y ​ ×100=23.2 %

Walen:

got it from toppr.com